OpenStudy (iwanttogotostanford):

What is the simplified form of the quantity of x plus 5, all over the quantity of 3x

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo Ive really learned a lot more from your help! I get it a lot easier

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo where'd you go?? I get my math a lot better with your teaching and help!

OpenStudy (anonymous):

@iwanttogotostanford I think you can handle this one. You've already walked through two examples that are similar to this question.

OpenStudy (iwanttogotostanford):

no i really can't I need help with just this last one! @Cardinal_Carlo

OpenStudy (iwanttogotostanford):

no more after this i can do it myself after

OpenStudy (iwanttogotostanford):

i stil have 20 questions like this and i just need one more example

OpenStudy (anonymous):

I know that you already know how to do this one, and neither I nor anyone will be holding your hand when you go to Stanford. You got this @iwanttogotostanford

OpenStudy (iwanttogotostanford):

i know its just that i have 20 more like this one but ok thanks!

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo I've tried it and i got the quantity of 4x squared plus 24x plus 31, all over the quantity of 4x plus 7

OpenStudy (anonymous):

Could you please use the equation function of OpenStudy, so that we can clearly see your expressions?

OpenStudy (iwanttogotostanford):

im not sure how to exactly @Cardinal_Carlo

OpenStudy (iwanttogotostanford):

I always end up messing it up

OpenStudy (anonymous):

Maybe you should get more practice with the Equation button first. It really helps. Otherwise, people have a hard time communicating answers with you. And yes, you may mess up a couple of times. I know I did the first few times I used it. But you know, you get used to it, and it becomes second-nature. It really is worth your while. Please get good at it. It makes it easier for the both of us.

OpenStudy (iwanttogotostanford):

ok @Cardinal_Carlo

OpenStudy (anonymous):

Ha ha I just noticed how this question is under a Spanish section. How funny.

OpenStudy (iwanttogotostanford):

oops wrong i didn't mean to shesh

OpenStudy (iwanttogotostanford):

4x2+24x+31/4x+7

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (anonymous):

Okay, good try. Does your expression actually look like this? \[\frac{ 4x ^{2} + 24x + 31 }{ 4x + 7 }\]

OpenStudy (iwanttogotostanford):

yes

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (anonymous):

From our starting expression: \[\frac{ x + 5 }{ 3x + 4 } + \frac{ x + 4 }{ x + 3 }\] we get: \[\frac{ x + 5 \left( x + 3 \right) + \left( x + 4 \right)\left( 3x + 4 \right) }{ 3x + 4 \left( x + 3 \right) }\] \[\frac{ x ^{2} + 8x + 15 + 3x ^{2} + 16x + 16 }{ 3x ^{2} + 13x + 12 }\] which simplifies to: \[\frac{ 4x ^{2} + 24x + 31 }{ 3x ^{2} + 13x + 12 }\] Is this one of the choices? @iwanttogotostanford

OpenStudy (iwanttogotostanford):

yes

OpenStudy (anonymous):

This is not spanish>>>????